after senior high you are decided to push thru your education in college. since it is difficult to be working student, when you are studying in college, you wanted to start saving now while working part time, because it is much easier due to online class. if you plan to deposit savings from your work now, which is Php.20,000 and the interest rate of your bank offers 5% compounded quarterly, represent the amount of your money in (t)time.
solve the word problem of exponential function and graph it
thank you so much in advance ms,mr

Answer by Theo(13342)

the bank is offering 5% compounded quarterly.
you deposit 20,000 now.
your exponential formula is y = a * b^x
a = 20,000
b = .10/4 = .025 + 1 = 1.025
formula becomes y = 20,000 * 1.025 ^ x
this formula can be graphed as shown below:


you can see that the formula is exponential from the long term view shown below.


if you let y = f and a = p and b = (1 + r) and x = n, then you get the formula:

f = p * (1 + r) ^ n

f is the future vlaue
p is the present value
r is the interest rate per time period
n is the number of time periods.

when p = 20,000 and r = 10% per year, the formula becomes:

f = 20,000 * (1 + .10/4) ^ n

.10/4 is the interest rate per quarter.
n is the number of quarters.

simplify this formula to get:

f = 20,000 * 1.025 ^ n

that's the same exponential formula as y = 20,000 * 1.025 ^ x, where a = 20,000 and b = 1.025